What mass of solid Lanthanum (III) oxalate nonahydrate { La2 ( C2O4 )3 ∙ 9 H2O } can be obtained from 650 mL of 0.0170 M aqueous solution of LaCl3 by adding a stoichiometric amount of sodium oxalate ?​

1
Expert's answer
2022-05-05T04:08:03-0400

n(LaCl3) = C(LaCl3) * V(LaCl3) = 0.0170 M * 0.650 L = 0.011 moles

m(LaCl3) = n(LaCl3) * Mr(LaCl3) = 0.011 moles * 245,26 g/mol = 2.698 g


2LaCl3(aq) + 3Na2C2O4(s) + 9H2O(l)  →  La2(C2O4)3·9H2O(s) + 6NaCl(aq)


m( La2(C2O4)3·9H2O) = (m(LaCl3) * Mr( La2(C2O4)3·9H2O)) / 2Mr(LaCl3)

m( La2(C2O4)3·9H2O) = (2.698 g * 703.87 g/mol) / 2 * 245,26 g/mol = 3.87 g